OpenStudy (anonymous):

HOW DO YOU DO THIS QUESTION? Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2.

OpenStudy (anonymous):

First off, what will the slope of a line that is parallel to y=x+2 be?

OpenStudy (anonymous):

they have the same slope

OpenStudy (anonymous):

Good, do you know how to use the point slope formula?

OpenStudy (anonymous):

\[y-y _{1} = m(x-x _{1})\]

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

Y-(-2)=1(X-(-4))

OpenStudy (anonymous):

Cool, now solve for y.

OpenStudy (anonymous):

the equation for the question is y=1/4x +2

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

Same principle. You know the slope will be the same. You've got a point on the line. Just plug and chug. :)

OpenStudy (anonymous):

That makes more sense cause the other one would give you the same equation

OpenStudy (anonymous):

The trickiest part of these is generally making sure that you change the signs as needed.

OpenStudy (anonymous):

Yeah, I noticed that. I thought it might have been a trick question.

OpenStudy (anonymous):

is it Y-(-2)=1/4 (X-(-4)?

OpenStudy (anonymous):

Yup. Now just solve for y.

OpenStudy (anonymous):

okk thanks alot

OpenStudy (anonymous):

No problem. Please not that doing this for a line that was perpendicular is almost exactly the same, you've just got to find the slope of the new line. Remember when lines are perpendicular: \[m _{1} \times m _{2} = -1\]

OpenStudy (anonymous):

is it y = 1/4x - 3

OpenStudy (anonymous):

I'm getting y = 1/4x - 1

OpenStudy (anonymous):

?????:(

OpenStudy (anonymous):

Here's our point slope formula again: \[y-(-2)=1/4 (x-(-4)\] Distribute the 1/4: \[y+2=1/4x +1\] Subtract 2 from both side: \[y=1/4x -1\]

OpenStudy (anonymous):

Gotta watch those sign changes :)

OpenStudy (anonymous):

oh ok i get it yea i messed up with the signs but thanks alot :))))))))))))))))))))))))

OpenStudy (anonymous):

Not a problem. :D Good luck!

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. is the answer to this y = -x +2??

OpenStudy (anonymous):

The slope of a line perpendicular to y = -x + 6 is m=1. To get that, take the original slope, flip it upside down, and reverse the sign.

OpenStudy (anonymous):

oh yea i forgot about that

OpenStudy (anonymous):

I've got to run, but just to make sure you've got it. m=1 \[y-(-2)=1(x-(-4))\] \[y+2=x+4\] y=x+2